### v1

@@ I Introduction
In ALCC a non-perturbative definition of ${\rho_{\rm vac}}$ was proposed based on thermodynamics. In the “Thermal” channel euclidean time is compactified on a circle of circumference $\beta=1/T$ where $T$ is the temperature. Thus states are the usual scattering states in infinite $d=D-1$ dimensional volume. This is in contrast to the spatially compactified channel “SpC”. where states partially live on a circle of one compactified spatial dimension. We showed that these two channels lead to the same result for free massive theories, however the Thermal channel is better behaved since most integrals are already convergent. The results of this paper are implicitly based on the Thermal channel in the limit $\beta\to 0$. Let us summarize the non-perturbative definition of ${\rho_{\rm vac}}$ in the thermal channel proposed in Mingling ; ALCC . Let $p$ denote the pressure and ${\cal F}$ the free energy density. Assuming the theory has a single mass scale $m$, such as the mass of the lightest particle, we define the scaling variable $r=m\beta$. Then one can write
@@ III.2 A prescription for ${\rho_{\rm vac}}$
A related but slightly different limit constrains the one-point functions. Even for the stress-energy tensor, these one-point form factors are constants and are not necessarily zero. For instance for the magnetic perturbation of the Ising model, $\Theta$ is the spin field $\sigma$ and has non-zero 1-particle form factors DelfinoMussardo ; DelfinoCardy ; DelfinoE8 . It was conjectured in these works that when a subset of the rapidities are taken to infinity, the form factors should factorize as a result of a kind of cluster decomposition. Let us assume this conjecture. In (33) if one only takes the $\theta_{1}\to\infty$ limit without also the subsequent limit $\theta_{1}\to\theta_{2}$ the form factors should factorize into 1-point form factors. We thus propose:
@@ III.4 Analytical details for the sinh-Gordon model
Form factors for the sinh-Gordon model were studied in great detail by Fring and Mussardo FringMussardo . Henceforth $F(\theta_{1},\theta_{2})=F(\theta_{12})$ is the two particle form factor for $\Theta$. From the Watson equation we define a minimal solution $F_{\rm min}$:
@@ IV.2 D=2 case in terms of Mandelstam variable $s$
$m^{2}=-i\lim_{s\to\infty|s=4m^{2}}\,F(4m^{2}-s)=\lim_{s\to\infty|s=m^{2}}\left(1-S(s)\right)\,\langle 0|\Theta|0\rangle.$ (77)
@@ IV.3 Prescription for ${\rho_{\rm vac}}$ in D=4 spacetime dimensions
$-i\lim_{s\to\infty|s=4m^{2}}\,F(4m^{2}-s)=-m^{2}\left(\lim_{s\to\infty|s=m^{2}}{\cal T}(s)\right)\,\langle 0|\Theta|0\rangle.$ (82)
@@
$\lim_{s\to\infty}{\cal T}(s)=\dfrac{i}{m^{2}}\,\dfrac{\mathfrak{g}}{s}.$ (83)
@@
$\langle 0|\Theta|0\rangle=\dfrac{1}{3}\,\dfrac{m^{4}}{\mathfrak{g}}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\rho_{\rm vac}}=\dfrac{1}{12}\,\dfrac{m^{4}}{\mathfrak{g}}.$ (84)
@@ V Potential applications to Beyond the Standard Model physics
${\rho_{\rm vac}}\approx\dfrac{1}{12}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}},$ (87)

### v2

@@ I Introduction
In ALCC a non-perturbative definition of ${\rho_{\rm vac}}$ was proposed based on thermodynamics. In the “Thermal” channel euclidean time is compactified on a circle of circumference $\beta=1/T$ where $T$ is the temperature. Thus states are the usual scattering states in infinite $d=D-1$ dimensional volume. This is in contrast to the spatially compactified channel “SpC”, where states partially live on a circle of one compactified spatial dimension. We showed that these two channels lead to the same result for free massive theories, however the Thermal channel is better behaved since most integrals are already convergent. The results of this paper are implicitly based on the Thermal channel in the limit $\beta\to 0$. Let us summarize the non-perturbative definition of ${\rho_{\rm vac}}$ in the thermal channel proposed in Mingling ; ALCC . Let $p$ denote the pressure and ${\cal F}$ the free energy density. Assuming the theory has a single mass scale $m$, such as the mass of the lightest particle, we define the scaling variable $r=m\beta$. Then one can write
@@ III.2 A prescription for ${\rho_{\rm vac}}$
A related but slightly different limit constrains the one-particle form factors. Even for the stress-energy tensor, these one-particle form factors are constants and are not necessarily zero. For instance for the magnetic perturbation of the Ising model, $\Theta$ is the spin field $\sigma$ and has non-zero 1-particle form factors DelfinoMussardo ; DelfinoCardy ; DelfinoE8 . It was conjectured in these works that when a subset of the rapidities are taken to infinity, the form factors should factorize as a result of a kind of cluster decomposition. Let us assume this conjecture. In (33) if one only takes the $\theta_{1}\to\infty$ limit without also the subsequent limit $\theta_{1}\to\theta_{2}$ the form factors should factorize into 1-particle form factors. We thus propose:
@@ III.4 Analytical details for the sinh-Gordon model
Form factors for the sinh-Gordon model were studied in great detail by Fring, Mussardo and Simonetti FringMussardo . Henceforth $F(\theta_{1},\theta_{2})=F(\theta_{12})$ is the two particle form factor for $\Theta$. From the Watson equation we define a minimal solution $F_{\rm min}$:
@@ IV.2 D=2 case in terms of Mandelstam variable $s$
$-i\,m^{2}=-i\lim_{s\to\infty|s=4m^{2}}\,F(4m^{2}-s)=\lim_{s\to\infty|s=m^{2}}\left(1-S(s)\right)\,\langle 0|\Theta|0\rangle.$ (77)
@@ IV.3 Prescription for ${\rho_{\rm vac}}$ in D=4 spacetime dimensions
$\lim_{s\to\infty|s=4m^{2}}\,F(4m^{2}-s)=m^{2}\left(\lim_{s\to\infty|s=m^{2}}{\cal T}(s)\right)\,\langle 0|\Theta|0\rangle.$ (82)
@@
$\lim_{s\to\infty}{\cal T}(s)=\dfrac{1}{m^{2}}\,\dfrac{\mathfrak{g}}{s}.$ (83)
@@
$\langle 0|\Theta|0\rangle=\,\dfrac{3m^{4}}{\mathfrak{g}}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\rho_{\rm vac}}=\dfrac{3}{4}\,\dfrac{m^{4}}{\mathfrak{g}}.$ (84)
@@ V Potential applications to Beyond the Standard Model physics
${\rho_{\rm vac}}\approx\dfrac{3}{4}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}},$ (87)